Grade 99 grade maths

Without actually calculating the cubes find the value of (-1)3​+(-2)3​+(-3)​3​+(-4)3​+​2(5)3​. Also say the identity used.

sushil
9 Years agoGrade 9

3 Answers

Arun
9 Years ago
(-1)^3​+(-2)^3​+(-3)^​3​+(-4)^3​+​2(5)^3
[(-1)^3​+(-4)^3​+​(5)^3​] + [(-2)^3​+(-3)​^3​+(5)^3​]
we know that when a+ b+ c = 0 then a^3 +b^3 +c^3 =  3* a*b*c
3[( –1)( –4)(5)] + 3[( –2)( –3)(5)]
60 + 30
 =150    answer
hope it helps
Arun
9 Years ago
sorry it was
60 + 90 = 150 (answer)
i wrote that by mistake.
hope you understand
you can do more question on similair pattern
rasheed
8 Years ago
we know that a+b+c = 0 then a3+b3+c3 = 3abc
 
(-1)3 + (- 4)3 + (5)3 = 3(-1)(-4)(5)  = 60
(-2)3 + (-3)3 + (5)3   = 3(-2)(-3)(5) = 90
= 60 + 90 = 150